# Diameter and mixing time of the giant component in the percolated hypercube

### v2

@@ 1. Introduction
The study of the percolated hypercube $Q^{d}_{p}$ was pioneered by Burtin [B77] and Sapoženko [S67] who showed that $p=1/2$ is a sharp threshold for connectivity of $Q^{d}_{p}$. A more drastic change of behaviour of typical instances of the model happens around $p=1/d$. Erdős and Spencer [ES79] observed that when $p=c/d$ with $c<1$, $Q^{d}_{p}$ typically contains only components of order $O(d)$, and Ajtai, Komlós, and Szemerédi [AKS81] completed the picture by showing that $Q^{d}_{p}$ typically contains a giant connected component $L_{1}$ with linearly many vertices when $c>1$. A subsequent strengthening of these results by Bollobás, Kohayakawa and Łuczak [BKL92] showed that, when $c>1$, the number of vertices in $L_{1}$ is typically $y2^{d}+o(2^{d})$ where $y:=y(c)$ is the survival probability of a Galton-Watson process with offspring distribution Poisson($c$), and typically all other (smaller) components have order $O(d)$, mirroring a similar phenomenon in the Erdős-Rényi random graph $G(n,p)$ [ER60].
@@
Despite the similarities, many properties of the giant component of the random graph $G(n,c/n)$ with $c>1$ were better understood thanks to the homogeneous nature of the host graph $K_{n}$. Two key examples are the typical diameter of the giant component and the asymptotic mixing time of a simple random walk on the giant component. The typical diameter of the giant component of $G(n,c/n)$ with $c>1$ was determined very precisely in a line of research [CL01, FR07, RW10] and turns out to be of asymptotic order $\Theta_{c}(\log n)$. The mixing time of a simple random walk, roughly speaking, measures the time needed by a simple random walk to forget its starting point (for a formal definition, see Section 6). The asymptotic mixing time of a simple random walk on the giant component in $G(n,c/n)$ with $c>1$ is of order $\Theta_{c}((\log n)^{2})$ [BKW14, FR08].
@@
The remarkable similarity between the Erdős-Rényi random graph and the percolated hypercube with the same (constant) average degree raises the question whether the typical diameter and the typical mixing time of a simple random walk on the giant component $L_{1}$ in $Q^{d}_{p}$ have the same quantitative behaviour. Already in 1994, Bollobás, Kohayakawa, and Łuczak [BKL94d, Problem 15] asked about the order of the typical diameter of $L_{1}$ and, in particular, if it is polynomial in $d$. The latter question was answered in a breakthrough of Erde, Kang, and Krivelevich [EKK22] who showed – using expansion properties of the hypercube – that the typical diameter of $L_{1}$ is of order $O(d^{3})$. This result was later improved by Diskin, Erde, Kang, and Krivelevich [DEKK24] to $O(d(\log d)^{2})$. In each of [DEKK24, EKK22] it was asked if, similarly to the giant of $G(n,c/n)$, $\Theta_{c}(d)$ is the right dependency on $d=\log_{2}|V(Q^{d})|$.
@@
In a separate line of research, Benjamini and Mossel [BM03, Section 3] asked whether the asymptotic mixing time of a (lazy) simple random walk on $L_{1}$ is of order $\Theta_{c}(d^{2})$, as in the case of the giant component of $G(n,c/n)$. This question was later reiterated by van der Hofstad and Nachmias [HN17, Open Problem (6)], and is advertised on the webpage of the Levin–Peres–Wilmer book [LPW17] “Markov Chains and Mixing Times” [Pereswebsite, Question 3]. Progress towards this conjecture was made by Erde, Kang and Krivelevich [EKK22] who provided an upper bound of $O(d^{11})$, and by Diskin, Erde, Kang and Krivelevich [DEKK24] who showed an upper bound of $O(d^{2}(\log d)^{2})$.
@@
Note that the lower bounds on each of the typical diameter and the typical mixing time follow from the existence of bare paths (that is, paths containing only vertices of degree 2) of length $\Omega_{c}(d)$ in $L_{1}$, which can be verified by a simple second moment computation (see, e.g., [EKK22, Section 5]).
@@
Let us note that the previous papers [DEKK24, EKK22] that achieved partial progress towards the conjectures resolved in Theorem 1 essentially relied on a ‘direct’ sprinkling argument and on the product structure of the hypercube. A key missing ingredient in the previous works was a tight expansion estimate, in particular for large connected sets, in the giant component. Naively, one could hope to obtain such an estimate using a ‘reverse’ sprinkling, also known as ‘thinning’, a process of removing edges with a given probability. Indeed, if a large set in the giant component has a weak expansion, removing each of the edges in its boundary is ‘fairly cheap’ and results in two large components — a very atypical structure. However, this does not immediately imply the probability bound of Theorem 2, as the set of deleted edges depends on the structure of the exposed random subgraph. It turns out that this constitutes the main challenge in resolving the conjectures. In this paper, we perform a quantitative analysis of small perturbations on the edges of the giant component, allowing us to resolve this challenge (see Section 2 for more details). Importantly, our arguments do not rely on the hypercube’s product structure, and so they could adapt to other sparse high-dimensional percolation models and to settings where conductance of large sets governs mixing behaviour (see, e.g., [AKLP23, AKP21, BBY08, CS24, JKM24, MD18, MMJ24, OCo98]).
As mentioned above, a key ingredient in the proof of Theorem 1(b) lies in Proposition 6.2 and Corollary 6.3, which describe the expansion properties of the giant. The following theorem is an abbreviated version of these two results, which resolves [EKK22, Question 5.1] and might be of independent interest.
@@
Since by the above (not necessarily connected) sets of linear (in $2^{d}$) size typically edge-expand by a factor of $\Omega(1/d)$, using techniques from [K19] (in particular, [K19, Lemma 2.6 and Theorem 2.7]) implies the existence of an $\Omega(1/d)$-expander in $L_{1}$ of linear order, resolving [EKK22, Question 5.1]. Indeed, the latter result is tight (up to a constant factor), as already noted in [EKK22, Claim 5.2].
Finally, complementing results on critical percolation [B-RBN24, BCVSS06, DKLP11, HS04, HN20] by pinning down the mixing time of a lazy simple random walk on the giant component in the barely-supercritical regime $dp=1+o(1)$ is an intriguing and natural next step, and we believe our techniques can provide key insights to it.
@@ 2. Outline of the proofs
Section 5 is dedicated to the proof of the lower tail in Theorem 2. Before delving into the proof’s outline, we first recall the broad strategy used in [AKS81, BKL92, K23] to show that whp the number of vertices in the giant component $L_{1}$ of $Q^{d}_{p}$ is at least $(1-o(1))y2^{d}$ (with $p,y$ as defined in Theorem 2). We note that a similar overall strategy was employed in [DEKK24, EKK22] to establish expansion properties of the giant component $L_{1}$. The argument goes through sprinkling, also known as multi-stage exposure. In the paragraph below, we give a brief sketch of that proof.
@@
We move to the proof of Theorem 1(b). In Section 6.1, we use the results in Section 5 to show that ‘large’ connected sets in $L_{1}\subseteq G_{2}$ typically expand well (Propositions 6.1 and 6.2). This shows Corollary 6.3, which also implies Theorem 3. Utilising expansion estimates to obtain bounds on the mixing time was used in previous works on hypercube percolation [DEKK24, EKK22]: a missing ingredient there was a good quantitative estimate on the expansion of ‘large’ sets, which is a major contribution of this paper. The proof of Proposition 6.1 is based on the following argument: assuming that a ‘large’ set $S$ in $G_{2}$ has a small edge-boundary, the probability that $S$ is disjoint from the largest component of $G_{1}$ (or itself becomes such) is significant. As it turns out, the probability bound on the event that $S$ is disjoint from the largest component in $G_{1}$ obtained from analysis of the sprinkling between $G_{1}$ and $G_{2}$ is tight in general. However, for our needs, we derive a stronger probability bound which can hold only for connected sets (see Section 6.1 and the discussion therein). To that end, we utilise the rather general stability principle (Lemma 5.6), which was also used to prove Theorem 2. Roughly speaking, Lemma 5.6 shows that it is very unlikely for a ‘large’ connected set in $G_{2}$ to disintegrate into mostly small components in $G_{1}$ (and, in particular, such a set is far from a type-1 component). Thus, with a significant probability, we must have many vertices in ‘big’ components in $G_{1}$ outside $L_{1}$ — an event whose probability we already estimate in Section 5. Combining the expansion properties of connected subsets in the giant component with previously known expansion estimates for $Q^{d}_{p}$ [DEKK24], one can derive Theorem 1(b) by using a theorem of Fountoulakis and Reed [FR07a] (stated here as 6.4). The latter result relates the mixing time of a lazy simple random walk on a graph to the expansion properties of connected sets therein.
@@ 3.2. Concentration inequalities
We start by presenting a version of the well-known Chernoff’s bound for binomial random variables (see, for example, [JLR00, Theorem 2.1]).
@@
Next, we state a variant of the bounded difference inequality (see, e.g., Theorem 3.9 in [M98] and Corollary 6 in [War16]).
@@
We end this section with a switching lemma reminiscent of [Wor99, Theorem 2.19]. Fix integers $k,d\geq 1$ and denote by $\mathcal{S}_{d}(k)$ the family of sequences of length $kd$ where every element in $[d]$ appears $k$ times. A switching consists of exchanging the positions of two elements in a sequence. For two sequences $\sigma_{1},\sigma_{2}\in\mathcal{S}_{d}(k)$, we write $\sigma_{1}\sim\sigma_{2}$ if they differ by a single switching.
@@
The lemma follows by combining (1) and Azuma’s inequality (see e.g. [JLR00, Theorem 2.25]).
@@ 3.3. Auxiliary graph-theoretic results
The following decomposition lemma appears as Lemma 2.1 in [ADILS25]. For graphs $H\subseteq G$ and a function $w:V(G)\to\mathbb{R}$, we write $w(H)=\sum_{v\in V(H)}w(v)$.
@@
Next, we state an estimate on the number of subtrees of a $d$-regular graph with fixed order and rooted in a particular vertex (see, e.g., [BFM98, Lemma 2] and [K98, Chapter 7]).
@@
On the one hand, (2) is obtained by multiplying $d^{k}$ and the coefficient of $(f(z))^{\ell}$ in front of $z^{k}$. On the other hand, by [CGHJK96, (2.36)],
@@
Next, we state the celebrated Harper’s edge-isoperimetric inequality for hypercubes, see [H64] and also [B67, H76, L64].
@@ 3.4. Auxiliary results on percolation
We start this section by stating the well-known Harris’ inequality, also known as the FKG inequality, see [AS16, Theorem 6.3.3]. For a finite set $\Lambda$, we say that a family $\mathcal{A}$ of subsets of $\Lambda$ is increasing if, for all subsets $A\subseteq B$ of $\Lambda$ with $A\in\mathcal{A}$, we also have that $B\in\mathcal{A}$. Denote by $\Lambda_{p}$ a random subset of $\Lambda$ containing every element independently and with probability $p$.
@@
Next, we present several lemmas on percolated hypercubes. The first one deals with matchings in $Q^{d}_{p}$ and appears as [EKK22, Lemma 2.9].
@@
We also make use of some expansion properties of connected subsets in the giant component established more generally in [DEKK24, Theorem 4], see also the remark in the beginning of page 747 therein.
@@ 3.5. Branching processes and the BFS algorithm
We begin with a lemma quantifying the extinction probability of a supercritical Galton-Watson process with large total progeny. Recall that, for a probability distribution $\mu$ on non-negative integers, a Galton-Watson process with offspring distribution $\mu$ is a branching process starting from a root where every vertex produces a random number of children with distribution $\mu$ independently of other vertices. The following lemma is a simplified version of [vdH18, Theorem 3.8].
@@
It is a classic fact (see e.g. [vdH18, Theorem 3.1]) that the extinction probability of a Galton-Watson process with distribution $\mu$ is given by the smallest solution in $[0,1]$ of the equation
@@
In particular, $\bar{y}(c):=1-y(c)$ is the smallest solution in $[0,1]$ to $f(c,q)=q$ where $f(c,q):=\mathrm{e}^{c(q-1)}$ while $\bar{b}(d,c/d):=1-b(d,c/d)$ is the smallest solution in $[0,1]$ to $f_{d}(c,q)=q$ where $f_{d}(c,q):=(1-c/d+cq/d)^{d}$. To establish the first equality in the statement of the lemma, we use the implicit function theorem (see e.g. [Spi65, Theorem 2-12]) to show that $\bar{y}$ is a differentiable function on the interval $(1,\infty)$. To this end, it suffices to verify that

### v3

@@ 1. Introduction
The study of the percolated hypercube $Q^{d}_{p}$ was pioneered by Burtin [16] and Sapoženko [53] who showed that $p=1/2$ is a sharp threshold for connectivity of $Q^{d}_{p}$. A more drastic change of behaviour of typical instances of the model happens around $p=1/d$. Erdős and Spencer [27] observed that when $p=c/d$ with $c<1$, $Q^{d}_{p}$ typically contains only components of order $O(d)$, and Ajtai, Komlós, and Szemerédi [1] completed the picture by showing that $Q^{d}_{p}$ typically contains a giant connected component $L_{1}$ with linearly many vertices when $c>1$. A subsequent strengthening of these results by Bollobás, Kohayakawa and Łuczak [13] showed that, when $c>1$, the number of vertices in $L_{1}$ is typically $y2^{d}+o(2^{d})$ where $y:=y(c)$ is the survival probability of a Galton-Watson process with offspring distribution Poisson($c$), and typically all other (smaller) components have order $O(d)$, mirroring a similar phenomenon in the Erdős-Rényi random graph $G(n,p)$ [26].
@@
Despite the similarities, many properties of the giant component of the random graph $G(n,c/n)$ with $c>1$ were better understood thanks to the homogeneous nature of the host graph $K_{n}$. Two key examples are the typical diameter of the giant component and the asymptotic mixing time of a simple random walk on the giant component. The typical diameter of the giant component of $G(n,c/n)$ with $c>1$ was determined very precisely in a line of research [17, 28, 52] and turns out to be of asymptotic order $\Theta_{c}(\log n)$. The mixing time of a simple random walk, roughly speaking, measures the time needed by a simple random walk to forget its starting point (for a formal definition, see Section 6). The asymptotic mixing time of a simple random walk on the giant component in $G(n,c/n)$ with $c>1$ is of order $\Theta_{c}((\log n)^{2})$ [8, 30].
@@
The remarkable similarity between the Erdős-Rényi random graph and the percolated hypercube with the same (constant) average degree raises the question whether the typical diameter and the typical mixing time of a simple random walk on the giant component $L_{1}$ in $Q^{d}_{p}$ have the same quantitative behaviour. Already in 1994, Bollobás, Kohayakawa, and Łuczak [14, Problem 15] asked about the order of the typical diameter of $L_{1}$ and, in particular, if it is polynomial in $d$. The latter question was answered in a breakthrough of Erde, Kang, and Krivelevich [25] who showed – using expansion properties of the hypercube – that the typical diameter of $L_{1}$ is of order $O(d^{3})$. This result was later improved by Diskin, Erde, Kang, and Krivelevich [21] to $O(d(\log d)^{2})$. In each of [21, 25] it was asked if, similarly to the giant of $G(n,c/n)$, $\Theta_{c}(d)$ is the right dependency on $d=\log_{2}|V(Q^{d})|$.
@@
In a separate line of research, Benjamini and Mossel [9, Section 3] asked whether the asymptotic mixing time of a (lazy) simple random walk on $L_{1}$ is of order $\Theta_{c}(d^{2})$, as in the case of the giant component of $G(n,c/n)$. This question was later reiterated by van der Hofstad and Nachmias [35, Open Problem (6)], and is advertised on the webpage of the Levin–Peres–Wilmer book [45] “Markov Chains and Mixing Times” [50, Question 3]. Progress towards this conjecture was made by Erde, Kang and Krivelevich [25] who provided an upper bound of $O(d^{11})$, and by Diskin, Erde, Kang and Krivelevich [21] who showed an upper bound of $O(d^{2}(\log d)^{2})$.
@@
Note that the lower bounds on each of the typical diameter and the typical mixing time follow from the existence of bare paths (that is, paths containing only vertices of degree 2) of length $\Omega_{c}(d)$ in $L_{1}$, which can be verified by a simple second moment computation (see, e.g., [25, Section 5]).
@@
Let us note that the previous papers [21, 25] that achieved partial progress towards the conjectures resolved in Theorem 1 essentially relied on a ‘direct’ sprinkling argument and on the product structure of the hypercube. A key missing ingredient in the previous works was a tight expansion estimate, in particular for large connected sets, in the giant component. Naively, one could hope to obtain such an estimate using a ‘reverse’ sprinkling, also known as ‘thinning’, a process of removing edges with a given probability. Indeed, if a large set in the giant component has a weak expansion, removing each of the edges in its boundary is ‘fairly cheap’ and results in two large components — a very atypical structure. However, this does not immediately imply the probability bound of Theorem 2, as the set of deleted edges depends on the structure of the exposed random subgraph. It turns out that this constitutes the main challenge in resolving the conjectures. In this paper, we perform a quantitative analysis of small perturbations on the edges of the giant component, allowing us to resolve this challenge (see Section 2 for more details). Importantly, our arguments do not rely on the hypercube’s product structure, and so they could adapt to other sparse high-dimensional percolation models and to settings where conductance of large sets governs mixing behaviour (see, e.g., [5, 6, 7, 18, 39, 40, 41, 48]).
As mentioned above, a key ingredient in the proof of Theorem 1(b) lies in Proposition 6.2 and Corollary 6.3, which describe the expansion properties of the giant. The following theorem is an abbreviated version of these two results, which resolves [25, Question 5.1] and might be of independent interest.
@@
Since by the above (not necessarily connected) sets of linear (in $2^{d}$) size typically edge-expand by a factor of $\Omega(1/d)$, using techniques from [43] (in particular, [43, Lemma 2.6 and Theorem 2.7]) implies the existence of an $\Omega(1/d)$-expander in $L_{1}$ of linear order, resolving [25, Question 5.1]. Indeed, the latter result is tight (up to a constant factor), as already noted in [25, Claim 5.2].
Finally, complementing results on critical percolation [12, 15, 20, 36, 37] by pinning down the mixing time of a lazy simple random walk on the giant component in the barely-supercritical regime $dp=1+o(1)$ is an intriguing and natural next step, and we believe our techniques can provide key insights to it.
@@ 2. Outline of the proofs
Section 5 is dedicated to the proof of the lower tail in Theorem 2. Before delving into the proof’s outline, we first recall the broad strategy used in [1, 13, 44] to show that whp the number of vertices in the giant component $L_{1}$ of $Q^{d}_{p}$ is at least $(1-o(1))y2^{d}$ (with $p,y$ as defined in Theorem 2). We note that a similar overall strategy was employed in [21, 25] to establish expansion properties of the giant component $L_{1}$. The argument goes through sprinkling, also known as multi-stage exposure. In the paragraph below, we give a brief sketch of that proof.
@@
We move to the proof of Theorem 1(b). In Section 6.1, we use the results in Section 5 to show that ‘large’ connected sets in $L_{1}\subseteq G_{2}$ typically expand well (Propositions 6.1 and 6.2). This shows Corollary 6.3, which also implies Theorem 3. Utilising expansion estimates to obtain bounds on the mixing time was used in previous works on hypercube percolation [21, 25]: a missing ingredient there was a good quantitative estimate on the expansion of ‘large’ sets, which is a major contribution of this paper. The proof of Proposition 6.1 is based on the following argument: assuming that a ‘large’ set $S$ in $G_{2}$ has a small edge-boundary, the probability that $S$ is disjoint from the largest component of $G_{1}$ (or itself becomes such) is significant. As it turns out, the probability bound on the event that $S$ is disjoint from the largest component in $G_{1}$ obtained from analysis of the sprinkling between $G_{1}$ and $G_{2}$ is tight in general. However, for our needs, we derive a stronger probability bound which can hold only for connected sets (see Section 6.1 and the discussion therein). To that end, we utilise the rather general stability principle (Lemma 5.6), which was also used to prove Theorem 2. Roughly speaking, Lemma 5.6 shows that it is very unlikely for a ‘large’ connected set in $G_{2}$ to disintegrate into mostly small components in $G_{1}$ (and, in particular, such a set is far from a type-1 component). Thus, with a significant probability, we must have many vertices in ‘big’ components in $G_{1}$ outside $L_{1}$ — an event whose probability we already estimate in Section 5. Combining the expansion properties of connected subsets in the giant component with previously known expansion estimates for $Q^{d}_{p}$ [21], one can derive Theorem 1(b) by using a theorem of Fountoulakis and Reed [29] (stated here as 6.4). The latter result relates the mixing time of a lazy simple random walk on a graph to the expansion properties of connected sets therein.
@@ 3.2. Concentration inequalities
We start by presenting a version of the well-known Chernoff’s bound for binomial random variables (see, for example, [38, Theorem 2.1]).
@@
Next, we state a variant of the bounded difference inequality (see, e.g., Theorem 3.9 in [47] and Corollary 6 in [55]).
@@
We end this section with a switching lemma reminiscent of [56, Theorem 2.19]. Fix integers $k,d\geq 1$ and denote by $\mathcal{S}_{d}(k)$ the family of sequences of length $kd$ where every element in $[d]$ appears $k$ times. A switching consists of exchanging the positions of two elements in a sequence. For two sequences $\sigma_{1},\sigma_{2}\in\mathcal{S}_{d}(k)$, we write $\sigma_{1}\sim\sigma_{2}$ if they differ by a single switching.
@@
The lemma follows by combining (1) and Azuma’s inequality (see e.g. [38, Theorem 2.25]).
@@ 3.3. Auxiliary graph-theoretic results
The following decomposition lemma appears as Lemma 2.1 in [4]. For graphs $H\subseteq G$ and a function $w:V(G)\to\mathbb{R}$, we write $w(H)=\sum_{v\in V(H)}w(v)$.
@@
Next, we state an estimate on the number of subtrees of a $d$-regular graph with fixed order and rooted in a particular vertex (see, e.g., [11, Lemma 2] and [42, Chapter 7]).
@@
On the one hand, (2) is obtained by multiplying $d^{k}$ and the coefficient of $(f(z))^{\ell}$ in front of $z^{k}$. On the other hand, by [19, (2.36)],
@@
Next, we state the celebrated Harper’s edge-isoperimetric inequality for hypercubes, see [32] and also [10, 33, 46].
@@ 3.4. Auxiliary results on percolation
We start this section by stating the well-known Harris’ inequality, also known as the FKG inequality, see [2, Theorem 6.3.3]. For a finite set $\Lambda$, we say that a family $\mathcal{A}$ of subsets of $\Lambda$ is increasing if, for all subsets $A\subseteq B$ of $\Lambda$ with $A\in\mathcal{A}$, we also have that $B\in\mathcal{A}$. Denote by $\Lambda_{p}$ a random subset of $\Lambda$ containing every element independently and with probability $p$.
@@
Next, we present several lemmas on percolated hypercubes. The first one deals with matchings in $Q^{d}_{p}$ and appears as [25, Lemma 2.9].
@@
We also make use of some expansion properties of connected subsets in the giant component established more generally in [21, Theorem 4], see also the remark in the beginning of page 747 therein.
@@ 3.5. Branching processes and the BFS algorithm
We begin with a lemma quantifying the extinction probability of a supercritical Galton-Watson process with large total progeny. Recall that, for a probability distribution $\mu$ on non-negative integers, a Galton-Watson process with offspring distribution $\mu$ is a branching process starting from a root where every vertex produces a random number of children with distribution $\mu$ independently of other vertices. The following lemma is a simplified version of [34, Theorem 3.8].
@@
It is a classic fact (see e.g. [34, Theorem 3.1]) that the extinction probability of a Galton-Watson process with distribution $\mu$ is given by the smallest solution in $[0,1]$ of the equation
@@
In particular, $\bar{y}(c):=1-y(c)$ is the smallest solution in $[0,1]$ to $f(c,q)=q$ where $f(c,q):=\mathrm{e}^{c(q-1)}$ while $\bar{b}(d,c/d):=1-b(d,c/d)$ is the smallest solution in $[0,1]$ to $f_{d}(c,q)=q$ where $f_{d}(c,q):=(1-c/d+cq/d)^{d}$. To establish the first equality in the statement of the lemma, we use the implicit function theorem (see e.g. [54, Theorem 2-12]) to show that $\bar{y}$ is a differentiable function on the interval $(1,\infty)$. To this end, it suffices to verify that